### v3

@@ 2.2.1 Detection of electron events
The final measurement with the PLATON prototype aimed to detect, for the first time ever with a plenoptic camera, single elementary particles in plastic scintillator on an event-by-event basis. A block of EJ-262 plastic scintillator [66] (parameters are listed in Sec. 4.1) was exposed to a source, placed on the top side. The electrons exiting the container of the source can carry an energy up to a maximum of 1.5 MeV. In order to maximise the photon collection, we had to position the scintillator block about 5 cm from the main lens of the PLATON prototype, thus out of focus. To certify the reconstruction of the electron position, two different datasets were collected, with and without the on top of the scintillator block. We name the two datasets, respectively, as and background. To decrease the probability of dark counts, the PLATON prototype was placed in a thermal chamber at a temperature of C, reducing the median dark count rate (DCR) to 0.002 cps/m2 (0.4 cps/pixel). The background sample allows us to identify noisy pixels, i.e. with a DCR higher than 0.5 cps/pixel, which were masked at the post-processing stage, further reducing the SwissSPAD2 photodetection efficiency (PDE) down to 2.5%, however increasing the signal-to-noise ratio. Thus, given the limited light yield of the prototype (f/2.4, active area reduced by the glue), only frames with at least four counts were selected, as these are more likely to contain at least two scintillation photons, necessary to fit a point in the object space. In the post-processing (see Sec. 4.2), candidate photons were identified if the relative minimum distance was less than 3 mm. Out of 9 frames with four counts found in the sample, only four events were selected upon the convergence of the post-processing. As shown in Fig. 3d, for each event the reconstructed position of was closer than 20 mm to its true position, dominated by the depth resolution, as discussed in Sec. 2.2. This result is consistent within the statistical uncertainty with the spatial resolution obtained for a low number of photons shown in Fig. 2. The same analysis was done on both the and the background sample with frames containing 3 counts. The rejection rates were, respectively, 61% and 81%. In the latter two cases, the distribution of the distance from the nominal position is evenly distributed across the entire depth range. Event displays of the four candidates can be found in Fig. 7.
2.3 Simulated particle detection with the PLATON detector
@@
As a physics study case, we chose the detection of GeV muon neutrinos. In fact, PLATON could be suitable for the near detector of future accelerator long-baseline neutrino oscillation (LBL) experiments, which will search for leptonic charge-parity violation and determine the neutrino mass ordering [1, 2, 3]. In these experiments, the precise measurement of the neutrino-nucleus cross section is crucial for the reduction of the key systematic uncertainties, and it relies on the high-resolution tracking and calorimetry of final-state low-momentum hadrons (e.g., multiple protons down to 200 MeV/c). We studied the performance of a small PLATON module consisting of two optical arrays, each of four plenoptic cameras, pointing to two orthogonal faces of a block of EJ-262 plastic scintillator (parameters are listed in Sec. 4.1). Performing the analysis described below would be challenging for a larger detector volume with the computing power currently available. Since the attenuation length is way longer than the scintillator volume, a perfect transparency was assumed. This assumption is justified even for larger volumes due to the availability of organic liquid scintillators with analogous light output but attenuation lengths of the order of 20 m or greater [67, 62]. The simulated detector is shown in Fig. 4a.
@@
Owing to the results obtained with the PLATON prototype (see Sec. 2.1), we designed the PLATON detector unit (a single camera). The adopted strategy consisted of optimising the lateral resolution, with a depth resolution of 1 mm over a depth of field of 10 cm. Exploiting the depth resolution to match the two orthogonal views, the result is a 3D spatial resolution that mainly belongs to the sub-millimetre lateral one, especially in relatively low-multiplicity events such as neutrino interactions. Both the MLA and the main lens were simulated with f/1. The diameter of the main lens is 50 mm, while that of the micro-lenses is 1 mm. Realistic parameters of an optimised SPAD array sensor, which we refer to as PLATON SPAD, were simulated. The PDP (40% peak at 500 nm), simulated as in Sec. 4.1 and the pixel pitch of with a 60% geometrical fill factor amount to a PDE of 24%. We neglect the tiling between adjacent photosensors. The SPAD time structure shown in Fig. 4b was simulated. Each count is measured with a 200 ps resolution timestamp. R&D on PLATON SPAD is ongoing to achieve the parameters above [68].
@@ 2.3.1 Single-point spatial resolution of an array of multiple cameras
Before studying the capability of a PLATON module to detect neutrinos, the 3D spatial resolution for a point light source equivalent to 1 MeV energy deposition was evaluated. The 3D spatial resolution was found to be 0.3 mm. For completeness, we compared plenoptic cameras with classical cameras, obtained by removing the MLA and adapting the lens-sensor distance, focusing the camera on the centre of the scintillator. While a single classical camera cannot precisely reconstruct the 3D position of the light source, we found that the resolution provided by multiple classical cameras is about four times worse than that of plenoptic cameras. On the other hand, we note that the size of the scintillator block is not that large and does not require a very long depth of field (DoF), making the performance of an array of classical cameras not too far from that of plenoptic cameras. As we will see later in Sec. 2.4, the resolution of a system of classical cameras deteriorates compared to plenoptic cameras for a m3 volume. Thus, we focus on the PLATON configuration with plenoptic cameras, as described in the previous section. For more details, we refer to Sec. 4.4.1.
2.3.2 Neutrino detection in PLATON
@@
The neutrino flux of the T2K LBL experiment [69, 70] was simulated parallel to both sensor planes, and uniformly in the scintillator volume with the NEUT 5.5.0 event generator [71]. Assuming that an external detector, such as a time projection chamber in a magnetised volume, would be able to detect and identify the muon escaping the scintillating volume, we simulated a sample of charged-current (CC) neutrino interactions, disregarding the contribution of neutral-current (NC) events.
@@
We developed a novel neural network based on the transformer architecture, commonly used in generative large language models, to efficiently capture the hyper-dimensional correlation between the detected scintillation photons. Details on the neural network are given in Sec. 4.5. We evaluated the particle tracking performance by computing the distance between each point (Geant4 output, see Sec. 4.1) along the truth trajectory of the particle and the closest reconstructed point (Figs. 4d and 4e). Overall, an average 3D-tracking resolution of approximately is obtained, with a range of to . When considering the distance between each reconstructed point and the closest truth point, the resolution becomes more consistent, with a narrower range of to , though the average remains unchanged. The resolution is smaller than m for events with three particles or fewer. When particle ranges are less than 3 mm, the resolution averages m. Overall, the results demonstrate that a tracking accuracy well below 1 mm is achievable within a sizeable PLATON module. A pattern recognition algorithm was developed to group the photon origins into different clusters that are then assigned to different particles. The results of the pattern recognition are shown in Fig. 5. The interaction vertex resolution for the full sample of CC interactions is 0.42 mm. The average difference between the number of identified and true particles is 0.25 with a standard deviation of 0.74. More details about the implementation of pattern recognition and particle identification can be found in Sec. 4.6.2.
@@
To demonstrate the capability of a PLATON detector module to detect neutrinos, we focus on the selection of both CC and CC events, noting that the latter are notoriously more challenging to identify due to the presence of higher multiplicity. From Fig. 5f, one can see that the momentum of stopping protons reconstructed from the range is inferred with a resolution better than 10% across almost the entire range, approximately 5% at the energy flux peak, and close to 10% only below 200 MeV/c. The CC selection purity is 90% up to 1.4 GeV/c momentum of the most energetic (leading) proton. The small background contamination comes mostly from events with higher proton multiplicity, typically produced with lower momenta. The total selection efficiency is 78%. It exceeds 80% for leading proton momenta above 300 MeV/, but drops below 50% for momenta above 1.1 GeV/, where the muon and the proton become more collinear. The proton reconstruction efficiency, shown for the CC sample, exhibits a 50% threshold at approximately 215 MeV/c. For the case of CC events, both the selection efficiency and the purity reach 92%. The direction of single protons is determined with an angular resolution of .
@@ 2.4 Single-point spatial resolution at 1 m3 scale
Finally, the spatial resolution for a point light source in a PLATON detector with a organic scintillator has been studied. Due to limited computational resources, we were unable to perform a neutrino simulation study similar to the one in Sec. 2.3.2, but we leave it for future work. The same analysis as in Sec. 2.3.1 was performed. As shown in Fig. 6, a 3D spatial resolution of 3.7 mm can be achieved for an energy loss of MeV (10,000 photons generated). For reference, this resolution corresponds to an effective segmentation of approximately 7.5 mm. Unlike a segmented detector, it improves up to 1.5 mm for an energy loss of MeV (100,000 photons generated), with an effective segmentation of 3 mm.
@@
With a few improvements, a sub-millimetre spatial resolution in a scintillator could be achieved. First, the optical parameters have not been optimised for such a volume but have been kept the same as for the module. Moreover, the used post-processing method adopts a simplified parametrisation of the MLA. On the other hand, we also observed that the NN described in Sec. 4.5 allowed for an additional improvement in spatial resolution. Thus, one might expect a similar improvement also in the tonne-scale PLATON. It is interesting to note that a smaller pixel pitch, though technically challenging, would allow for better constraint of the light field. Preliminary studies showed 30% improvement in spatial resolution for a 1 m pitch without any adaptation of the optics.
@@ 3 Discussion
In this work, we propose a change of paradigm in the high-resolution detection of elementary particles in dense scintillator-based detectors. The key is the use of imaging photosensors with single-photon sub-nanosecond time resolution, such as SPAD array sensors, integrated into a system of plenoptic cameras to provide unambiguous 3D images of elementary particles interacting within a monolithic volume of scintillator. We referred to this novel detector concept as PLATON.
We built and successfully characterised the spatial resolution of the PLATON detector prototype, the first plenoptic camera ever instrumented with a SPAD array sensor, and we developed a novel post-processing method suitable for photon-starved images. We also succeeded in reconstructing the position of 90Sr electrons on an event-by-event basis. All the results are in agreement with our simulations, which gives us confidence in designing a realistic PLATON detector for neutrino detection.
Building on the experience gained through prototyping, we carried out a simulation-based analysis of muon neutrino interactions in a PLATON detector, employing an improved yet realistic design. We developed a new high-performance image post-processing based on deep learning and complemented by pattern recognition algorithms. A tracking resolution down to 200 m, along with high purities and efficiencies in the selection of neutrino interactions with multi-proton final states, was achieved.
@@
The PLATON detector configuration offers broader applicability beyond accelerator-based neutrino studies, which were adopted here as a case study. Its high spatial resolution makes this concept well-suited for neutrinoless double-beta decay (-less ) experiments such as Ref. [72], and potentially even for tonne-scale detectors (see Ref. [73] for a comprehensive review). Furthermore, following the design philosophy outlined in Ref. [47], the PLATON concept appears promising for application in very-large-scale neutrino detectors [74, 67, 75, 76]. Similarly, it may enhance the reconstruction of light rings characteristic of Cherenkov radiation [77, 78].
To conclude, our work paves the way towards high-resolution particle tracking and calorimetry in unsegmented volumes. The development of SPAD array imaging sensors, or other photosensor technologies with analogous features, is the key. We are currently conducting R&D to develop a SPAD array sensor that meets the requirements discussed in this article, specifically a PDE similar to that of high-dynamic range SiPMs [31] and single-photon sub-nanosecond time resolution.
The PLATON concept therefore opens up new areas of application for deploying plenoptic cameras for imaging, including positron-emission tomography (PET), neutron radiography, muon tomography, synthetic computed tomography (CT), and proton CT.

### v4

@@ 2.2.1 Detection of electron events
The final measurement with the PLATON-prototype aimed to detect, with a plenoptic camera, single elementary particles in plastic scintillator on an event-by-event basis. A block of EJ-262 plastic scintillator [datasheet_EJ_260_262] (parameters are listed in Sec. 4.1) was exposed to a source, placed on the top side. The electrons exiting the container of the source can carry an energy up to a maximum of 1.5 MeV. In order to maximise the photon collection, we had to position the scintillator block about 5 cm from the main lens of the PLATON-prototype, just outside the optimal depth of field (150 mm), and hence out of focus. This is limiting the angular resolution on the photon direction. To certify the reconstruction of the electron position, two different datasets were collected, with and without the on top of the scintillator block. We name the two datasets, respectively, as and background. To decrease the probability of dark counts, the PLATON-prototype was placed in a thermal chamber at a temperature of C, reducing the median dark count rate (DCR) to 0.002 cps/m2 (0.4 cps/pixel). The background sample allows us to identify noisy pixels, i.e. with a DCR higher than 0.5 cps/pixel, which were masked at the post-processing stage, further reducing the SwissSPAD2 photodetection efficiency (PDE) down to 2.5%, however increasing the signal-to-noise ratio. Thus, given the limited light yield of the prototype (f/2.4, active area reduced by the glue), only frames with at least four counts were selected, as these are more likely to contain at least two scintillation photons, necessary to fit a point in the object space. In the post-processing (see Sec. 4.2), candidate photons were identified if the relative minimum distance was less than 3 mm. Out of 9 frames with four counts found in the sample, only four events were selected upon the convergence of the post-processing. As shown in Fig. 3d, for each event the reconstructed position of was closer than 20 mm to its true position, dominated by the depth resolution, as discussed in Sec. 2.2.
This result is consistent, within statistical uncertainty, with the spatial resolution obtained for a low number of photons when accounting for the distance between the main lens and the scintillator, which is not in focus, as shown in Fig. 2.
The same analysis was done on both the and the background sample with frames containing 3 counts. The rejection rates were, respectively, 61% and 81%. In the latter two cases, the distribution of the distance from the nominal position is evenly distributed across the entire depth range. Event displays of the four candidates can be found in Fig. 7. From a toy Monte Carlo simulation of the background sample we expect 0.17 background events with 4 or more counts to be reconstructed between 0 and 20 mm from the true position.
2.3 Simulated particle detection with the PLATON-10cm detector
@@
As a physics study case, we chose the detection of GeV muon neutrinos. In fact, the PLATON concept could be suitable for the near detector of future accelerator long-baseline neutrino oscillation (LBL) experiments, which will search for leptonic charge-parity violation and determine the neutrino mass ordering [Hyper-Kamiokande:2025fci, Hyper-Kamiokande:2018ofw, Abi:2020qib]. In these experiments, the precise measurement of the neutrino-nucleus cross section is crucial for the reduction of the key systematic uncertainties, and it relies on the high-resolution tracking and calorimetry of final-state low-momentum hadrons (e.g., multiple protons down to 200 MeV/c). We studied the performance of a small PLATON module consisting of two optical arrays, each of four plenoptic cameras, pointing to two orthogonal faces of a block of EJ-262 plastic scintillator (parameters are listed in Sec. 4.1), which we call PLATON-10cm. Performing the analysis described below would be challenging for a larger detector volume with the computing power currently available. Since the attenuation length is way longer than the scintillator volume, a perfect transparency was assumed. This assumption is justified even for larger volumes due to the availability of organic liquid scintillators with analogous light output but attenuation lengths of the order of 20 m or greater [JUNO:2021vlw, ABUSLEME2021164823]. The simulated detector is shown in Fig. 4a.
@@
Owing to the results obtained with the PLATON-10cm prototype (see Sec. 2.1), we designed the PLATON-10cm detector unit (a single camera). The adopted strategy consisted of optimising the lateral resolution, with a depth resolution of 1 mm over a depth of field of 10 cm. Exploiting the depth resolution to match the two orthogonal views, the result is a 3D spatial resolution that mainly belongs to the sub-millimetre lateral one, especially in relatively low-multiplicity events such as neutrino interactions. Both the MLA and the main lens were simulated with f/1. The diameter of the main lens is 50 mm, while that of the micro-lenses is 1 mm. Realistic parameters of an optimised SPAD array sensor, which we refer to as PLATON SPAD, were simulated. The PDP (40% peak at 500 nm), simulated as in Sec. 4.1 and the pixel pitch of with a 60% geometrical fill factor amount to a PDE of 24%. We neglect the tiling between adjacent photosensors. With all efficiencies accounted for, 3,670 photons were detected per event. The SPAD time structure shown in Fig. 4b was simulated. Each count is measured with a 200 ps resolution timestamp. R&D on PLATON SPAD is ongoing to achieve the parameters above [Kaneyasu2025PlatonSPAD].
@@ 2.3.1 Single-point spatial resolution of an array of multiple cameras
Before studying the capability of a PLATON-10cm module to detect neutrinos, the 3D spatial resolution for a point light source equivalent to 1 MeV energy deposition was evaluated. Of the 10,000 photons produced, on average 141 photons were detected. The 3D spatial resolution was found to be 0.3 mm. For completeness, we compared plenoptic cameras with classical cameras, obtained by removing the MLA and adapting the lens-sensor distance, focusing the camera on the centre of the scintillator. Here an average of 149 photons were detected per event. While a single classical camera cannot precisely reconstruct the 3D position of the light source, we found that the resolution provided by multiple classical cameras is about four times worse than that of plenoptic cameras. Comparing a setup with two plenoptic cameras to a setup with two classical cameras, we see a significant difference in reconstruction efficiency, indicated in Fig. 8. While the two plenoptic cameras are able in all events to reconstruct the 3D origin of the point source, when the event only falls in the view of a single classical camera, the reconstruction fails, leading to around 23% of failed reconstruction attempts. On the other hand, we note that the size of the scintillator block is not that large and does not require a very long depth of field (DoF), making the performance of an array of classical cameras not too far from that of plenoptic cameras. As we will see later in Sec. 2.4, the resolution of a system of classical cameras deteriorates compared to plenoptic cameras for a m3 volume. Thus, we focus on the PLATON-10cm configuration with plenoptic cameras, as described in the previous section. For more details, we refer to Sec. 4.4.1.
2.3.2 Neutrino detection in PLATON-10cm
@@
The neutrino flux of the T2K LBL experiment [Abe_2015_Measurements_of_neutrino, FluxPrediction2016] was simulated parallel to both sensor planes, and uniformly in the scintillator volume with the NEUT 5.5.0 event generator [Hayato_2021_The_NEUT_neutrino]. Assuming that an external detector, such as a time projection chamber with a spatial resolution and angular resolution of 1 mm and 0.05 rad respectively in a magnetised volume, would be able to detect and identify the muon escaping the scintillating volume, we simulated a sample of charged-current (CC) neutrino interactions, disregarding the contribution of neutral-current (NC) events.
@@
We developed a custom neural network based on the transformer architecture, commonly used in generative large language models, to efficiently capture the hyper-dimensional correlation between the detected scintillation photons. Details on the neural network are given in Sec. 4.5. We evaluated the particle tracking performance by computing the distance between each point (Geant4 output, see Sec. 4.1) along the truth trajectory of the particle and the closest reconstructed point (Figs. 4d and 4e). Overall, an average 3D-tracking resolution of approximately is obtained, with a range of to . When considering the distance between each reconstructed point and the closest truth point, the resolution becomes more consistent, with a narrower range of to , though the average remains unchanged. The resolution is smaller than m for events with three particles or fewer. When particle ranges are less than 3 mm, the resolution averages m. Overall, the results demonstrate that a tracking accuracy well below 1 mm is achievable within a sizeable PLATON module. This is in agreement with the results of 4.4 and 2.3.1. A pattern recognition algorithm was developed to group the photon origins into different clusters that are then assigned to different particles. The results of the pattern recognition are shown in Fig. 5. The interaction vertex resolution for the full sample of CC interactions is 0.42 mm. The average difference between the number of identified and true particles is 0.25 with a standard deviation of 0.74. More details about the implementation of pattern recognition and particle identification can be found in Sec. 4.6.2.
@@
To demonstrate the capability of a PLATON-10cm detector module to detect neutrinos, we focus on the selection of both CC and CC events, noting that the latter are notoriously more challenging to identify due to the presence of higher multiplicity. From Fig. 5f, one can see that the momentum of stopping protons reconstructed from the range is inferred with a resolution better than 10% across almost the entire range, approximately 5% at the energy flux peak, and close to 10% only below 200 MeV/c. The CC selection purity is 90% up to 1.4 GeV/c momentum of the most energetic (leading) proton. The small background contamination comes mostly from events with higher proton multiplicity, typically produced with lower momenta. The total selection efficiency is 78%. It exceeds 80% for leading proton momenta above 300 MeV/, but drops below 50% for momenta above 1.1 GeV/, where the muon and the proton become more collinear. The proton reconstruction efficiency, shown for the CC sample, exhibits a 50% threshold at approximately 215 MeV/c. For the case of CC events, both the selection efficiency and the purity reach 92%. The direction of single protons is determined with an angular resolution of .
@@ 2.4 Single-point spatial resolution at 1 m3 scale
Finally, the spatial resolution for a point light source in a PLATON detector with a organic scintillator has been studied, named PLATON-1m. Due to limited computational resources, we were unable to perform a neutrino simulation study similar to the one in Sec. 2.3.2, but we leave it for future work. The same analysis as in Sec. 2.3.1 was performed. As shown in Fig. 6, a 3D spatial resolution of 3.7 mm can be achieved for an energy loss of MeV (10,000 photons generated). For reference, this resolution corresponds to an effective segmentation of approximately 7.5 mm. Unlike a segmented detector, it improves up to 1.5 mm for an energy loss of MeV (100,000 photons generated), with an effective segmentation of 3 mm.
@@
With a few improvements, a sub-millimetre spatial resolution in a scintillator could be achieved. First, the optical parameters have not been optimised for such a volume but have been kept the same as for the module. Moreover, the used post-processing method adopts a simplified parametrisation of the MLA. On the other hand, we also observed that the NN described in Sec. 4.5 allowed for an additional improvement in spatial resolution. Thus, one might expect a similar improvement also in the tonne-scale PLATON-1m. It is interesting to note that a smaller pixel pitch, though technically challenging, would allow for better constraint of the light field. Preliminary studies showed 30% improvement in spatial resolution for a 1 m pitch without any adaptation of the optics.
@@ 3 Discussion
In this work, we propose a change of paradigm in the high-resolution detection of elementary particles in dense scintillator-based detectors. The key is the use of imaging photosensors with single-photon sub-nanosecond time resolution, such as SPAD array sensors, integrated into a system of plenoptic cameras to provide unambiguous 3D images of elementary particles interacting within a monolithic volume of scintillator. We referred to this detector concept as PLATON.
We built and successfully characterised the spatial resolution of the PLATON detector prototype, a plenoptic camera instrumented with a SPAD array sensor, and we developed a post-processing method suitable for photon-starved images. We also succeeded in reconstructing the position of 90Sr electrons on an event-by-event basis. All the results are in agreement with our simulations, which gives us confidence in designing a realistic PLATON detector for neutrino detection.
Building on the experience gained through prototyping, we carried out a simulation-based analysis of muon neutrino interactions in a PLATON detector, employing an improved yet realistic design. We developed a high-performance image post-processing based on deep learning and complemented by pattern recognition algorithms. A tracking resolution down to 200 m, along with high purities and efficiencies in the selection of neutrino interactions with multi-proton final states, was achieved.
@@
The PLATON detector configuration offers broader applicability beyond accelerator-based neutrino studies, which were adopted here as a case study. Its calorimetric capability combined with high spatial resolution could make this concept potentially suitable for scintillator-based neutrino-less double-beta decay (-less ) experiments and makes it worth studying. Examples could be Ref. [SuperNEMO] or large-scale detectors such as Ref. [KamLAND-Zen:2024eml, SNO:2021xpa, AUTY2023168204, SNO:2020fhu] (see also Ref. [Dolinski:2019nrj] for a comprehensive review). Furthermore, following the design philosophy outlined in Ref. [Dalmasson:2017pow], the PLATON concept appears promising for application in very-large-scale neutrino detectors [DAYABAY, JUNO:2021vlw, SNO:2025qqy, Theia:2019non]. Similarly, it may enhance the reconstruction of light rings characteristic of Cherenkov radiation [Jiang:2019xwn, PEREGO2012606]. Compared to liquid argon time projection chambers (LArTPCs), currently the only technology capable of 3D tracking particles produced by neutrino interactions with O(mm) resolution in giant volumes, the PLATON technology would profit of its much faster response. Thus, it is ideal for the detection and the reconstruction of the time-of-flight of fast neutrons produced by neutrinos. All would be achieved without the need of big cryogenic infrastructures. On the other hand, its scalability to the kilotonne scale achieved by LArTPCs will have to be proven.
To conclude, our work paves the way towards high-resolution particle tracking and calorimetry in unsegmented volumes. The development of SPAD array imaging sensors, or other photosensor technologies with analogous features, is the key. We are currently conducting R&D to develop a SPAD array sensor that meets the requirements discussed in this article, specifically a PDE similar to that of high-dynamic range SiPMs [hamamatsu:mppc] and single-photon sub-nanosecond time resolution.
Owing to high-spatial resolution in large and dense active volumes combined with an excellent time resolution, the PLATON concept therefore naturally opens up new areas of application for deploying plenoptic cameras for imaging, including positron-emission tomography (PET), neutron radiography, muon tomography, synthetic computed tomography (CT), and proton CT.